P and q are two sets such that np=17, nq=14 and n(p n q)=5. find n(p u q).

Answer by ikleyn(52781)

.

n(p U q) = n(p) + n(q) - n(p ∩ q) = 17 + 14 - 5 = 26.    ANSWER

Solved and explained.